# QERA: an Analytical Framework for Quantization Error Reconstruction

### v1

@@ Abstract
The growing number of parameters and computational demands of large language models (LLMs) present significant challenges for their efficient deployment. Recently, there is an increasing interest in quantizing weights to extremely low precision while offsetting the resulting error with low-rank, high-precision error reconstruction terms. The combination of quantization and low-rank approximation is now popular in both adapter-based, parameter-efficient fine-tuning methods such as LoftQ [1] and low-precision inference techniques including ZeroQuant-V2 [2]. Usually, the low-rank terms are calculated via the singular value decomposition (SVD) of the weight quantization error, minimizing the Frobenius and spectral norms of the weight approximation error. Recent methods like LQ-LoRA [3] and LQER [4] introduced hand-crafted heuristics to minimize errors in layer outputs (activations) rather than weights, resulting improved quantization results.
However, these heuristic-based methods lack an analytical solution to guide the design of quantization error reconstruction terms. In this paper, we revisit this problem and formulate an analytical framework, named Quantization Error Reconstruction Analysis (QERA), and offer a closed-form solution to the problem. We show QERA benefits both existing low-precision fine-tuning and inference methods – QERA achieves a fine-tuned accuracy gain for $\Delta_{\text{acc}}$ = 6.05% of 2-bit RoBERTa-base on GLUE compared to LoftQ; and obtains $\Delta_{\text{acc}}$ = 2.97% higher post-training quantization accuracy of 4-bit Llama-3.1-70B comapred to ZeroQuant-V2 and $\Delta_{\text{ppl}}$ = -- 0.28 lower perplexity on WikiText2 compared to LQER.
Index Terms:
LLMs, low-rank approximation, parameter-efficient fine-tuning, post-training quantization
I Introduction
@@
The demand for efficient deployment of large language models (LLMs) has been increasing [5]. LLMs now typically contain billions of parameters [6, 7], making their fine-tuning and inference computationally expensive and resource-intensive [8]. To address these challenges, there has been a surge of interest in building efficient fine-tuning and inference methods. One popular formulation is to apply a low-rank term to reconstruct the error after quantization. Given a linear layer ${\bm{y}}={\bm{x}}{\bm{W}}$, the weight matrix ${\bm{W}}\in{\mathbb{R}}^{m\times n}$ is quantized to $\widetilde{{\bm{W}}}$, and we rewrite ${\bm{W}}\approx\widetilde{{\bm{W}}}+{\bm{A}}_{k}{\bm{B}}_{k}$ such that both ${\bm{A}}_{k}$ and ${\bm{B}}_{k}$ are low-rank yet high-precision terms with rank $k\ll\min(m,n)$.
@@
We call the problem of finding the optimal ${\bm{A}}_{k}$ and ${\bm{B}}_{k}$ quantization error reconstruction. Interestingly, this problem has, coincidentally, seen widespread application in two actively researched areas: quantized parameter-efficient fine-tuning (QPEFT) and post-training quantization (PTQ) for model inference. QPEFT refers to fine-tuning techniques that adapt LLMs to specific tasks by quantizing pretrained weights and updating only a small number of extra parameters, hence significantly reducing memory requirements and training time, such as QLoRA [3]. On the other side, PTQ is a training-free method that reduces the model size and may accelerate the forward pass if the underlying hardware supports it. Recently, researchers combined PTQ with quantization error reconstruction [2, 9, 4] to further reduce weight precision. Works such as ZeroQuant-V2 [2] and LQER [4] have shown that adding a high-precision low-rank component, as low as 8 or 32, can recover considerable model performance for 3- or 4-bit weight quantization.
@@
Although both the QPEFT and PTQ methods have demonstrated substantial performance improvements in lowering the computational overhead of LLMs, a theoretical analysis of quantization error reconstruction is lacking. Usually, ${\bm{A}}_{k}$ and ${\bm{B}}_{k}$ are calculated by applying truncated singular value decomposition (SVD) to the weight quantization error $({\bm{W}}-\widetilde{{\bm{W}}})$, minimizing the Frobenius and spectral norms of the weight approximation error. However, recent work on activation-aware quantization and knowledge distillation implies that minimizing layer output error generally leads to a greater performance gain than minimizing weight approximation error [10, 9, 11].
@@
Besides the unsettled minimization objective, it has remained unclear whether there exists a theoretically optimal solution for the values of ${\bm{A}}_{k}$ and ${\bm{B}}_{k}$, and if so, how one can solve for it. A better initialization or theoretically grounded initialization of ${\bm{A}}_{k}$ and ${\bm{B}}_{k}$ brings direct benefits for both QPEFT and PTQ. In QPEFT, the initialization of LoRA [12], which uses element-wise Gaussian random values for ${\bm{A}}_{k}$ and zeros for ${\bm{B}}_{k}$, struggles under aggressive quantization since the quantization error can derail fine-tuning. In PTQ, the quantized model performance is based on the computation of the low-rank terms, given a specific quantization function $\mathrm{q}(\cdot)$ and rank $k$.
@@ II Related Work
LoRA [12] is a representative PEFT method that introduces trainable low-rank terms to adapt the model to a specific task. Take a linear layer as an example,
@@
QLoRA [3] extends LoRA by quantizing the pretrained weights stored in GPU memory to reduce memory footprint.
@@
where $\mathrm{dq}(\cdot)$ is the dequantization function. QLoRA introduces weight quantization error $({\bm{W}}-\widetilde{{\bm{W}}})$, shifting the starting point of fine-tuning. To address this problem, LoftQ [1] initializes the adapter using the SVD-based low-rank approximation of $({\bm{W}}-\widetilde{{\bm{W}}})$ to reduce the weight approximation error:
@@
LQ-LoRA [3] also adopts LoftQ’s iterative method but keeps track of a scaled variant of the objective, $\operatorname*{arg\,min}_{{\bm{A}}_{k},{\bm{B}}_{k}}||{\bm{D}}_{\text{row}}({\bm{W}}-\widetilde{{\bm{W}}}-{\bm{A}}_{k}{\bm{B}}_{k}){\bm{D}}_{\text{col}}||_{F}$, where ${\bm{D}}_{\text{row}}$ and ${\bm{D}}_{\text{col}}$ are heuristic homogenous row/column matrices from activation statistics. LQ-LoRA exits the iteration when the scaled objective function stops decreasing due to the lack of a theoretical justification for LoftQ.
@@
ZeroQuant-V2 [2] is the earliest weight-only quantization method introducing low-rank quantization error reconstruction to the PTQ problem. They apply SVD to the weight quantization error $({\bm{W}}-\widetilde{{\bm{W}}})$ to calculate ${\bm{A}}_{k}$ and ${\bm{B}}_{k}$ (equivalent to LoftQ with one iteration). Combining low-rank terms and fine-grained quantization, ZeroQuant-V2 recovers the performance of 4-bit LLMs to a level comparable to 8-bit.
Recent quantization works have shown that activation statistics play a crucial role in weight-only LLM quantization [13, 10]. QLLM [9] trains the low-rank terms using gradient descent with a loss function that minimizes the output error of the attention layer. LQER [4] applies an activation-induced heuristic scale matrix ${\bm{S}}$ to the quantization error before calculating SVD, ${\bm{U}}{\bm{\Sigma}}{\bm{V}}^{T}=\mathrm{SVD}({\bm{S}}({\bm{W}}-\widetilde{{\bm{W}}}))$, and assigns ${\bm{A}}_{k}:={\bm{S}}^{-1}{\bm{U}}_{:,:k},\;{\bm{B}}_{k}:={\bm{\Sigma}}_{:k,:k}{\bm{V}}^{T}_{:k,:}$ (Refer to Algorithm 2 in the Appendix). LQER achieves significant improvement over ZeroQuant-V2 and observes that in some layers singular values are shaped toward a more desirable distribution where singular values decay faster. Note that ZeroQuant-V2 can also be considered as a special case where ${\bm{S}}$ in LQER is an identity matrix.
In summary, most of existing methods on QPEFT and PTQ target the minimization of the weight approximation error. Several recent works such as LQ-LoRA, QLLM, and LQER introduce activation-induced heuristics to the calculation of adapters/low-rank terms, but a justification for the optimization objective and the corresponding analytical framework are still missing.
III Our Analytical Framework
@@ III-A Problem Statement
Solution to Problem 1. From the Eckart-Young-Mirsky theorem [14], the optimal solution to Problem 1 with respect to rank $k$ is the truncated SVD of the weight error matrix:
@@
As noted in Section II, most existing works [1, 2, 3] in QPEFT and PTQ adopt this solution. However, we know that minimizing the weight approximation error is not equivalent to minimizing the layer output error. Furthermore, does minimizing the weight approximation error for each layer in a network effectively reduce the final model output error? We will show that the answer is negative in Section IV-B.
@@
Problem 2 motivates some recent works [9, 3, 4] to involve activation-induced heuristics in the optimization of ${\bm{C}}_{k}$. However, to our best knowledge, there is no work providing an analytical solution to this problem. In the following two sections, we will derive the analytical solution to Problem 2. More precisely, we present two solutions: one exact solution in Section III-B and an approximated solution based on a suitable statistical assumption in Section III-C.
III-B QERA-exact: Analytical Solution
QERA-exact is our exact solution to Problem 2. QERA-exact is computationally expensive as it calculates the autocorrelation matrix of the input space ${\mathbb{X}}$. However, as we will show in Section IV, QERA-exact recovers significant model performance in extremely low-precision quantization.
@@ III-B QERA-exact: Analytical Solution
${\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}^{\frac{1}{2}}$ is positive semi-definite. In the event that it has a zero eigenvalue, it would be normal to add a small diagonal perturbation to recover invertibility. In practice, we ran extensive experiments and find that ${\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}^{\frac{1}{2}}$ is invertible for all the pretrained models and datasets we present in Section IV.
@@
Using the property of the element-wise product [15], the RHS of the above can be simplified.
@@
Since ${\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}$ is a symmetric positive semi-definite matrix, it always has precisely one matrix square root, denoted as ${\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}^{\frac{1}{2}}$, that is also symmetric and positive semi-definite [16]. We reorganize Equation 15 as the following since both ${\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}$ and $({\bm{P}}{\bm{P}}^{T})$ are symmetric and positive semi-definite:
@@
$\operatorname*{arg\,min}_{{\bm{Q}}}||{\bm{Q}}_{k}-{\bm{Q}}||_{F}^{2}$ (18)
Note that multiplication by the invertible matrix ${\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}^{\frac{1}{2}}$ (Remark 1) does not change the rank of the matrix ${\bm{C}}_{k}$. According to the Eckart-Young-Mirsky theorem [14], the optimal rank $k$ approximation to ${\bm{Q}}_{k}$ is the truncated SVD of ${\bm{Q}}$:
@@
III-C QERA-approx: An Analytical Solution with the Uncorrelated Assumption
QERA-approx is our analytical solution to Problem 2 based on the assumption that different embedding dimensions are uncorrelated. This solution is more computationally efficient than the exact solution, and the assumption is testable on real-world datasets. The complete proof of QERA-approx is in Section A-B.
@@ III-C QERA-approx: An Analytical Solution with the Uncorrelated Assumption
We test this assumption on LLMs in Section V.
@@
For implementation, we assign ${\bm{A}}_{k}:={\bm{S}}^{-1}{\bm{U}}_{:,:k}$ and ${\bm{B}}_{k}:={\bm{\Sigma}}_{:k,:k}{\bm{V}}_{:k,:}$ to form the low-rank terms to save the memory and computation cost. Interestingly, QERA-approx solution is similar to the activation-induced heuristics in LQER [4], which calibrates the average absolute value on the embedding dimension (Refer to Algorithm 2 in the Appendix). In Section IV-C, we will show that our solution is more effective in practice and resolves the discrepancy between the recovered model performance and the number of calibration samples in LQER.
IV Experiments
In this section, we first introduce the experiment setup in Section IV-A. Then we present the results of our experiments on QPEFT and PTQ in Section IV-B and Section IV-C respectively.
IV-A Experiment Setup
We perform fine-tuning and quantization experiments separately, and compare with their respective SoTA methods. The experiments take around 6400 GPU hours in total. The hardware platform, separate GPU hours, software dependencies, and random seed settings can be found in Section A-C.
For fine-tuning experiments, we use Theorem 2, noted as QERA-approx, to initialize low-rank terms, and compare with full-finetuning, LoRA [12], QLoRA [17], and LoftQ [1]. Specifically, we adopt 5-iteration LoftQ, which is the officially recommended setup. We include both encoder-only model experiments (fine-tuning RoBERTa-base [18] on GLUE [19]) and decoder-only LLM experiments (fine-tuning LLaMA-2 [20] and LLaMA-3.1 [7] on continuous pretraining task SlimPajama [21] and supervised fine-tuning task GSM8K [22]). For each method/baseline, we sweep the learning rate and record the best result. The final results are averaged over three random seeds. The learning rate ranges and batch sizes are listed in Section A-C1.
For quantization experiments, we use both Theorem 1, noted as QERA-exact, and Theorem 2 (QERA-approx) to calculate the low-rank error reconstruction terms and report results separately. We compare with BF16, quantized model without error reconstruction terms ($w$-only), ZeroQuant-V2 [2], and LQER [4] at different precision setups. We also include HQQ [23], a leading 4-bit method that does not use quantization error reconstruction. We quantize LLMs of various sizes and model family, including TinyLlama [24], Gemma-2 [25], Phi-3.5 [26] and LLaMA-2/-3.1 [20, 7]. We use lm-evaluation-harness to report results on Wikitext2 [27], ARC (challenge) [28], BoolQ [29], CommonSenseQA [30], Winogrande [31], MMLU [32], and BigBench-Hard [33]. We also evaluate instruction-tuned model, Vicuna-v1.5 [34], with AlpacaEval 2.0 [35], which is an automatic evaluation tool for instruction-following tasks. Detailed setup is in Section A-C2.
IV-B Improved QPEFT

### v2

@@ Abstract
The growing number of parameters and computational demands of large language models (LLMs) present significant challenges for their efficient deployment. Recently, there is an increasing interest in quantizing weights to extremely low precision while offsetting the resulting error with low-rank, high-precision error reconstruction terms. The combination of quantization and low-rank approximation is now popular in both adapter-based, parameter-efficient fine-tuning methods such as LoftQ (Li et al., 2023) and low-precision inference techniques including ZeroQuant-V2 (Yao et al., 2023). Usually, the low-rank terms are calculated via the singular value decomposition (SVD) of the weight quantization error, minimizing the Frobenius and spectral norms of the weight approximation error. Recent methods like LQ-LoRA (Guo et al., 2023) and LQER (Zhang et al., 2024a) introduced hand-crafted heuristics to minimize errors in layer outputs (activations) rather than weights, resulting improved quantization results. However, these heuristic-based methods lack an analytical solution to guide the design of quantization error reconstruction terms. In this paper, we revisit this problem and formulate an analytical framework, named Quantization Error Reconstruction Analysis (QERA), and offer a closed-form solution to the problem. We show QERA benefits both existing low-precision fine-tuning and inference methods – QERA achieves a fine-tuned accuracy gain for $\Delta_{\text{acc}}$ = 6.05% of 2-bit RoBERTa-base on GLUE compared to LoftQ; and obtains $\Delta_{\text{acc}}$ = 2.97% higher post-training quantization accuracy of 4-bit Llama-3.1-70B compared to ZeroQuant-V2 and $\Delta_{\text{ppl}}$ = -- 0.28 lower perplexity on WikiText2 compared to LQER. We open-source our code and models at github.com/ChengZhang-98/QERA.
1 Introduction
@@
The demand for efficient deployment of large language models (LLMs) has been increasing (Faiz et al., 2023). LLMs now typically contain billions of parameters (Kaplan et al., 2020; Dubey et al., 2024), making their fine-tuning and inference computationally expensive and resource-intensive (Ding et al., 2023). To address these challenges, there has been a surge of interest in building efficient fine-tuning and inference methods. One popular formulation is to apply a low-rank term to reconstruct the error after quantization. Given a linear layer ${\bm{y}}={\bm{x}}{\bm{W}}$, the weight matrix ${\bm{W}}\in{\mathbb{R}}^{m\times n}$ is quantized to $\widetilde{{\bm{W}}}$, and we rewrite ${\bm{W}}\approx\widetilde{{\bm{W}}}+{\bm{A}}_{k}{\bm{B}}_{k}$ such that both ${\bm{A}}_{k}$ and ${\bm{B}}_{k}$ are low-rank yet high-precision terms with rank $k\ll\min(m,n)$.
@@
We call the problem of finding the optimal ${\bm{A}}_{k}$ and ${\bm{B}}_{k}$ quantization error reconstruction. Interestingly, this problem has, coincidentally, seen widespread application in two actively researched areas: quantized parameter-efficient fine-tuning (QPEFT) and post-training quantization (PTQ) for model inference. QPEFT refers to fine-tuning techniques that adapt LLMs to specific tasks by quantizing pretrained weights and updating only a small number of extra parameters, hence significantly reducing memory requirements and training time, such as QLoRA (Guo et al., 2023). On the other side, PTQ is a training-free method that reduces the model size and may accelerate the forward pass if the underlying hardware supports it. Recently, researchers combined PTQ with quantization error reconstruction (Yao et al., 2023; Liu et al., 2023a; Zhang et al., 2024a) to further reduce weight precision. Works such as ZeroQuant-V2 (Yao et al., 2023) and LQER (Zhang et al., 2024a) have shown that adding a high-precision low-rank component, as low as 8 or 32, can recover considerable model performance for 3- or 4-bit weight quantization.
@@
Although both the QPEFT and PTQ methods have demonstrated substantial performance improvements in lowering the computational overhead of LLMs, a theoretical analysis of quantization error reconstruction is lacking. Usually, ${\bm{A}}_{k}$ and ${\bm{B}}_{k}$ are calculated by applying truncated singular value decomposition (SVD) to the weight quantization error $({\bm{W}}-\widetilde{{\bm{W}}})$, minimizing the Frobenius and spectral norms of the weight approximation error. However, recent work on activation-aware quantization and knowledge distillation implies that minimizing layer output error may lead to a greater performance gain than minimizing weight approximation error (Lin et al., 2024; Liu et al., 2023a; Shao et al., 2023).
@@
Besides the unsettled minimization objective, it has remained unclear whether there exists a theoretically optimal solution for the values of ${\bm{A}}_{k}$ and ${\bm{B}}_{k}$, and if so, how one can solve for it. A better initialization or theoretically grounded initialization of ${\bm{A}}_{k}$ and ${\bm{B}}_{k}$ brings direct benefits for both QPEFT and PTQ. In QPEFT, the initialization of LoRA (Hu et al., 2021), which uses element-wise Gaussian random values for ${\bm{A}}_{k}$ and zeros for ${\bm{B}}_{k}$, struggles under aggressive quantization since the quantization error can derail fine-tuning. In PTQ, the quantized model performance is based on the computation of the low-rank terms, given a specific quantization function $\mathrm{q}(\cdot)$ and rank $k$.
@@ 2 Related Work
LoRA Hu et al. (2021) is a representative PEFT method that introduces trainable low-rank terms to adapt the model to a specific task. Take a linear layer as an example,
@@
QLoRA (Guo et al., 2023) extends LoRA by quantizing the pretrained weights stored in GPU memory to reduce memory footprint.
@@
where $\mathrm{dq}(\cdot)$ is the dequantization function. QLoRA introduces weight quantization error $({\bm{W}}-\widetilde{{\bm{W}}})$, shifting the starting point of fine-tuning. To address this problem, LoftQ (Li et al., 2023) initializes the adapter using the SVD-based low-rank approximation of $({\bm{W}}-\widetilde{{\bm{W}}})$ to reduce the weight approximation error:
@@
LQ-LoRA (Guo et al., 2023) also adopts LoftQ’s iterative method but keeps track of a scaled variant of the objective, $\operatorname*{arg\,min}_{{\bm{A}}_{k},{\bm{B}}_{k}}||{\bm{D}}_{\text{row}}({\bm{W}}-\widetilde{{\bm{W}}}-{\bm{A}}_{k}{\bm{B}}_{k}){\bm{D}}_{\text{col}}||_{F}$, where ${\bm{D}}_{\text{row}}$ and ${\bm{D}}_{\text{col}}$ are heuristic homogenous row/column matrices from activation statistics. LQ-LoRA exits the iteration when the scaled objective function stops decreasing due to the lack of a theoretical justification for LoftQ.
@@
ZeroQuant-V2 (Yao et al., 2023) is the earliest weight-only quantization method introducing low-rank quantization error reconstruction to the PTQ problem. They apply SVD to the weight quantization error $({\bm{W}}-\widetilde{{\bm{W}}})$ to calculate ${\bm{A}}_{k}$ and ${\bm{B}}_{k}$ (equivalent to LoftQ with one iteration). Combining low-rank terms and fine-grained quantization, ZeroQuant-V2 recovers the performance of 4-bit LLMs to a level comparable to 8-bit.
Recent quantization works have shown that activation statistics play a crucial role in weight-only LLM quantization (Liu et al., 2023b; Lin et al., 2024). QLLM (Liu et al., 2023a) trains the low-rank terms using gradient descent with a loss function that minimizes the output error of the attention layer. LQER (Zhang et al., 2024a) applies an activation-induced heuristic scale matrix ${\bm{S}}$ to the quantization error before calculating SVD, ${\bm{U}}{\bm{\Sigma}}{\bm{V}}^{T}=\mathrm{SVD}({\bm{S}}({\bm{W}}-\widetilde{{\bm{W}}}))$, and assigns ${\bm{A}}_{k}:={\bm{S}}^{-1}{\bm{U}}_{:,:k},\;{\bm{B}}_{k}:={\bm{\Sigma}}_{:k,:k}{\bm{V}}^{T}_{:k,:}$ (Refer to Algorithm 2 in the Appendix). LQER achieves significant improvement over ZeroQuant-V2 and observes that in some layers singular values are shaped toward a more desirable distribution where singular values decay faster. Note that ZeroQuant-V2 can also be considered as a special case where ${\bm{S}}$ in LQER is an identity matrix. To our knowledge, CALDERA (Saha et al., 2024) is the concurrent work close to ours. CALDERA focuses on a different problem setup to find optimal $\widetilde{W}$, $A_{k}$, $B_{k}$ all in low-precision formats that minimizes output error, with a lemma agreeing with our exact solution. We elaborate the connection and difference between CALDERA and QERA in Section A.3.
In summary, to solve the quantization error reconstruction (QER) problem, most of existing methods target the minimization of the weight approximation error. Several recent works such as LQ-LoRA, QLLM, and LQER introduce activation-induced heuristics to the calculation of adapters/low-rank terms, but a justification for the optimization objective and the corresponding analytical framework are still missing.
3 Our Analytical Framework
@@ 3.1 Problem Statement
Solution to Problem 1. From the Eckart-Young-Mirsky theorem (Eckart & Young, 1936), the optimal solution to Problem 1 with respect to rank $k$ is the truncated SVD of the weight error matrix:
@@
As noted in Section 2, most existing works (Li et al., 2023; Yao et al., 2023; Guo et al., 2023) in QPEFT and PTQ adopt this solution. However, we know that minimizing the weight approximation error is not equivalent to minimizing the layer output error. Furthermore, does minimizing the weight approximation error for each layer in a network effectively reduce the final model output error? We will show that the answer is negative in Section 4.2.
@@
Problem 2 motivates some recent works (Liu et al., 2023a; Guo et al., 2023; Zhang et al., 2024a) to involve activation-induced heuristics in the optimization of ${\bm{C}}_{k}$ but without a theoretical foundation. In the following two sections, we will derive the analytical solution to Problem 2. More precisely, we present two solutions: one exact solution in Section 3.2 and an approximated solution based on a suitable statistical assumption in Section 3.3.
3.2 QERA-exact: Analytical Solution
QERA-exact is our exact solution to Problem 2. QERA-exact is computationally expensive as it calculates the autocorrelation matrix of the input space ${\mathbb{X}}$. However, as we will show in Section 4, QERA-exact recovers significant model performance in extremely low-precision quantization.
@@ 3.2 QERA-exact: Analytical Solution
${\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}^{\frac{1}{2}}$ is positive semi-definite. In the event that it has a zero eigenvalue, it would be normal to add a small diagonal perturbation to recover invertibility. In practice, we ran extensive experiments and find that ${\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}^{\frac{1}{2}}$ is invertible for all the pretrained models and datasets we present in Section 4.
@@
Using the property of the element-wise product (Styan, 1973), the RHS of the above can be simplified.
@@
Since ${\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}$ is a symmetric positive semi-definite matrix, it always has precisely one matrix square root, denoted as ${\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}^{\frac{1}{2}}$, that is also symmetric and positive semi-definite (Horn & Johnson, 2012). We reorganize Equation 15 as the following since both ${\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}$ and $({\bm{P}}{\bm{P}}^{T})$ are symmetric and positive semi-definite:
@@
$\operatorname*{arg\,min}_{{\bm{Q}}_{k}}||{\bm{Q}}_{k}-{\bm{Q}}||_{F}^{2}$ (18)
Note that multiplication by the invertible matrix ${\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}^{\frac{1}{2}}$ (Remark 1) does not change the rank of the matrix ${\bm{C}}_{k}$. According to the Eckart-Young-Mirsky theorem (Eckart & Young, 1936), the optimal rank $k$ approximation to ${\bm{Q}}_{k}$ is the truncated SVD of ${\bm{Q}}$:
@@
3.3 QERA-approx: An Analytical Solution with the Uncorrelated Assumption
QERA-approx is our analytical solution to Problem 2 based on the assumption that different embedding dimensions are uncorrelated. This solution is more computationally efficient than the exact solution, and the assumption is testable on real-world datasets. The complete proof of QERA-approx is in Section A.2.
@@ 3.3 QERA-approx: An Analytical Solution with the Uncorrelated Assumption
We test this assumption on LLMs in Section 5.
@@
For implementation, we assign ${\bm{A}}_{k}:={\bm{S}}^{-1}{\bm{U}}_{:,:k}$ and ${\bm{B}}_{k}:={\bm{\Sigma}}_{:k,:k}{\bm{V}}_{:k,:}$ to form the low-rank terms to save the memory and computation cost. Interestingly, QERA-approx solution is similar to the activation-induced heuristics in LQER (Zhang et al., 2024a), which calibrates the average absolute value on the embedding dimension (Refer to Algorithm 2 in the Appendix). In Section 4.3, we will show that our solution is more effective in practice and resolves the discrepancy between the recovered model performance and the number of calibration samples in LQER.
4 Experiments
In this section, we first introduce the experiment setup in Section 4.1. Then we present the results of our experiments on QPEFT and PTQ in Section 4.2 and Section 4.3 respectively.
4.1 Experiment Setup
We perform QPEFT and PTQ experiments separately, and compare with their respective SoTA methods. The experiments take around 6400 GPU hours in total. The hardware platform, separate GPU hours, software dependencies, and random seed settings can be found in Section A.4.
For QPEFT experiments, we use Theorem 2, noted as QERA-approx, to initialize low-rank terms, and compare with full-finetuning, LoRA (Hu et al., 2021), QLoRA (Dettmers et al., 2024), and LoftQ (Li et al., 2023). Specifically, we adopt 5-iteration LoftQ, which is the officially recommended setup. We include both encoder-only model experiments (fine-tuning RoBERTa-base (Liu, 2019) on GLUE (Ye et al., 2019)) and decoder-only LLM experiments (fine-tuning LLaMA-2 (Touvron et al., 2023) and LLaMA-3.1 (Dubey et al., 2024) on continuous pretraining task SlimPajama (Soboleva et al., 2023) and supervised fine-tuning task GSM8K (Cobbe et al., 2021)). For each method/baseline, we sweep the learning rate and record the best result. The final results are averaged over three random seeds. The learning rate ranges and batch sizes are listed in Section A.4.1.
For PTQ experiments, we use both Theorem 1, noted as QERA-exact, and Theorem 2 (QERA-approx) to calculate the low-rank error reconstruction terms and report results separately. We compare with BF16, quantized model without error reconstruction terms ($w$-only), ZeroQuant-V2 (Yao et al., 2023), and LQER (Zhang et al., 2024a) at different precision setups. We also include HQQ (Badri & Shaji, 2023), a leading 4-bit method that does not use quantization error reconstruction. We quantize LLMs of various sizes and model family, including TinyLlama (Zhang et al., 2024b), Gemma-2 (Team et al., 2024), Phi-3.5 (Abdin et al., 2024) and LLaMA-2/-3.1 (Touvron et al., 2023; Dubey et al., 2024). We use lm-evaluation-harness to report results on Wikitext2 (Merity et al., 2016), ARC (challenge) (Clark et al., 2018), BoolQ (Clark et al., 2019), CommonSenseQA (Talmor et al., 2019), Winogrande (Sakaguchi et al., 2019), MMLU (Hendrycks et al., 2021), and BigBench-Hard (Suzgun et al., 2022). We also evaluate instruction-tuned model, Vicuna-v1.5 (Zheng et al., 2023), with AlpacaEval 2.0 (Dubois et al., 2024), which is an automatic evaluation tool for instruction-following tasks. Detailed setup is in Section A.4.2.
4.2 Improved QPEFT